# Improving Precision in Kinematic Weak Lensing with MIRoRS: Model-Independent Restoration of Reflection Symmetries

### v1

@@ 3.7 Outlier Rejection
Figure 5: Axis ratios calculated from Illustris. An overall uncertainty of 0.087 was found, with similar values across inclinations of $20^{\circ}$, $40^{\circ}$, and $60^{\circ}$.
4 Results and Discussion
4.1 Axis Errors
@@
We also analyzed our idealized mock data set, finding an error of $2.3^{\circ}$, in line with the findings of Häussler et al. (2007) for idealized galaxies. Given the idealized nature of this data, we would expect that features like bars, warps, and spirals would produce a larger uncertainty in a real observation. However, a good fit to the ideal data provides validation of the methods used for finding the position angle. The difference between the results from Illustris and idealized data gives an indication of the impact of non-ideal features, such as bars or warps, on the observed position angle.
Turning our attention to the errors in the axis ratio, shown in Figure 5, we find the Illustris sample is well-fit with a beta distribution across all inclination angles. The beta distribution is given by
$f(q;\alpha,\beta)=\frac{1}{B(\alpha,\beta)}q^{\alpha-1}(1-q)^{\beta-1},$ (56)
where $\alpha$ and $\beta$ are fitting parameters and $B(\alpha,\beta)$ is the beta function which is itself defined in terms of the gamma function,
$B(\alpha,\beta)=\frac{\Gamma(\alpha)\Gamma(\beta)}{\Gamma(\alpha+\beta)}.$ (57)
Furthermore, the variance can be computed from $\alpha$ and $\beta$ directly
$\sigma_{q}^{2}=\frac{\alpha\beta}{(\alpha+\beta)^{2}(\alpha+\beta+1)}.$ (58)
An overall uncertainty of 0.087 was found over the entirety of the dataset. That value is improved to between $\sigma_{q}=0.078$ and $\sigma_{q}=0.085$ if we consider each inclination individually. However, that information would not be available in a real observation, so we opt to use the larger value in our calculations. Note, the axis ratio does not match what we would expect considering an inclined thin disk. Instead, all axis ratios tend greater than a theoretical value, with discrepancy increasing with inclination. We believe this to be an effect due to limitations in the ideal thin disk calculation, rather than an error in the fitting. The axis ratio in the calculation of Equation 51 depends on the true observed ratio, not on a thin-disk approximation.
@@
Figure 6: Characterization of the algorithm across the mock data set. A fiducial set of parameters: $\gamma_{\times}=0$, $\phi=0^{\circ}$, $i=40^{\circ}$, and $\delta_{V}=10\text{ km s}^{-1}$, was used while the individual parameters were varied. Across each trial is 50 realizations of the same configuration. In the violin plots, orange lines mark the median of the data while the boxes give the interquartile range and whiskers extend to 5th and 95th percentiles. The combined model is given in purple, kinematic only in green, and a parametric kinematic model in blue.
4.2 Idealized Mock Data
@@
With the uncertainty in the photometric axes established, the model is then characterized on an idealized mock data set, described in Section 3.6.1. A set of fiducial values of $\gamma_{\times}=0$, $\phi=0^{\circ}$, $i=40^{\circ}$, and $\delta_{V}=10\text{ km s}^{-1}$ were held constant across all runs, with the exception of the parameter being studied. A total of 50 realizations of the data are used for each parameter configuration.
@@
We compared three approaches to fitting the data: (1) a nonparametric model using only kinematic data, (2) a nonparametric model using both kinematic and photometric data, and (3) a simple parametric model using only kinematic data based on Equation 53, henceforth referred to as kinematic, combined, and parametric respectively. The results are shown in Figure 6.
@@
While varying the value of $\gamma_{\times}$ from $-0.2$ to $0.2$, we find the kinematic approach generally outperforms the parametric model, with the addition of photometric data model improving over the kinematic only. The combined approach performs noticeably better at lower input shear. This is a possible indication on the limits of the approximations used in deriving the theoretical difference between the kinematic and photometric position angles, particularly in the small angle approximation used in Equation 13.
@@
The effects of inclination are most dramatic for the parametric model, with no ability to fit at the lowest inclination angle. The kinematic model performs better than the parametric at low inclination and improves with inclination up to $60^{\circ}$. The combined model performs well over a wide range of inclinations, with the smallest error at mid-inclinations from $40^{\circ}$ to $60^{\circ}$. Increasing to the highest inclination tested, $80^{\circ}$, all models have greater uncertainty than at mid-inclinations. Both nonparametric approaches give similar results at higher inclination.
@@
The information available from the kinematic and photometric data varies with inclination. The information available in the photometric data is greatest at low inclinations. The apparent axis ratio increases with decreasing inclination, causing Equation 20 to diverge as $q$ approaches unity and trend to zero as the galaxy approaches edge-on. The velocity field, however, is featureless at low inclinations. The line of sight velocity is a function of $\sin i$, with the differential velocity increasing with inclination. As the galaxy approaches edge-on, the minor axis vanishes in an infinitely thin disk model, causing the increase in uncertainty we see at $i=80^{\circ}$.
Figure 7: Impact of $\gamma_{+}$ on the fit for $\gamma_{\times}$ using the combined model on idealized mock data. The same fiducial set of parameters: $\gamma_{\times}=0$, $\phi=0^{\circ}$, $i=40^{\circ}$, and $\delta_{V}=10\text{ km s}^{-1}$, was used as in Figure 6. The upper-right panel shows the error in fitting $\gamma_{\times}$ with reference to a fixed detector coordinate system while the red lines indicate the true value of $\gamma_{\times}$ in a coordinate system aligned with the galaxy axes. The lower-right panel shows the effect of $\gamma_{+}$ on the fiducial set-up.
@@
The position angle shows dramatic effects as the angle approaches $45^{\circ}$. The kinematic only model is unable to fit for $\gamma_{\times}$ at $\phi=45^{\circ}$. After the algorithm fits for the position angle, application of $\gamma_{\times}$ shear does not have any effect on the symmetry. With the galaxy at $45^{\circ}$, $\gamma_{\times}$ (applied with reference to the coordinate axes) will behave as $\gamma_{+}$ and only produce stretching and compressing along the galaxy axes and have no effect on galaxy symmetry. The combined model, on the other hand, seems to strongly outperform all other models at that position angle. This is not the result of a good fit, but rather is due to the fitting process giving $\gamma_{\times}=0$ for any value of true shear. If we revisit Equation 25, we can see that at $\phi=45^{\circ}$, the detector frame kinematic position angle will also be $45^{\circ}$. This will push $\gamma_{\times}$ to zero in Equation 18, regardless of true shear. This is the greatest difference in the two different nonparametric approaches, with both approaches failing for different reasons. However, it could be possible to construct the methods in such a way that they are sensitive to the cross component of shear in reference to the galactic axes, rather than the prescribed coordinate system. This is left for future work.
@@
Finally, we look at the effect of noise added to the kinematic data. Gaussian noise is added at 1, 5, 10, and 20 km s-1 on the central pixel and scales with the inverse of the square root of the local intensity, given by Equation 54. All models show trends of a better fit at low noise and poor fit at high noise. The combined approach increasingly outperforms the other methods as noise increases. The quality of the parametric fit at extremely low-noise is probably an artifact of the mock data and parametric model using the same functional form and would not translate into performance in a real observation.
4.2.1 Effect of $\gamma_{+}$
@@
While our technique focuses primarily on solving for $\gamma_{\times}$, real observations will contain both components of shear. We explore the impact of $\gamma_{+}$ on our fit for $\gamma_{\times}$, again using idealized mock data as in the previous section. We do not attempt to fit for $\gamma_{+}$, but believe our methods could be improved by considering the parameter, which we leave for future work.
As seen in Figure 7, particularly the lower-right panel, $\gamma_{+}$ adversely affects the uncertainty in $\gamma_{\times}$ while having minimal impact on the median. The effects are more pronounced at both high values of $\gamma_{+}$ (lower-right panel) as well as $\gamma_{\times}$ (upper-left). Some slight bimodality at low inclinations could imply a systematic effect that vanishes when considering an average.
The upper-right panel of Figure 7 shows how position angle affects the fit for $\gamma_{\times}$ in the fixed detector coordinates. In these coordinates, a fiducial value of $\gamma_{\times}=0$ is applied with varied values for $\gamma_{+}$. As the position angle approaches $45^{\circ}$, $\gamma_{+}$ applied in the fixed coordinates act as $\gamma_{\times}$ in a coordinate system aligned with the galaxy axes. This value of $\gamma_{\times}$ in the galaxy coordinates is shown with a red line at each configuration.
While $\gamma_{+}$ does not change the symmetry of a galaxy, it does affect our ability to treat the transformation of the photometric axes as a simple rotation, thus the derivation done in §2.1 no longer captures the full picture. Furthermore, this indicates the problem cannot be resolved by the simple inclusion of $\gamma_{+}$ in our transformation matrix $\mathcal{A}$, which is only used for symmetry transformations of the velocity field.
@@
4.3 Illustris Data
Extending our analysis to the Illustris data, we fit the 358 halos at three inclinations each ($20^{\circ}$, $40^{\circ}$, and $60^{\circ}$) with both the kinematic and combined models. These halos are from a z=0 snapshot and all expected to have $\gamma_{\times}=0$. Results are shown in Table 3, along with a comparison to analysis of the same data using a parametric model by DW23. Comparing the two nonparametric models, shown in Figure 8, we see the combined model shows a distinct improvement over the kinematic model. Uncertainty in the kinematic model is $0.05$ considering all inclination angles and improves to $0.039$ at an inclination of $60^{\circ}$. The combined model has an uncertainty of $0.039$, improving to $0.027$ at $i=60^{\circ}$.
Figure 8: Comparison of the nonparametric kinematic and combined models. Adding photometric data provides a clear improvement on the fit, across all inclination angles.

### v2

@@ III.7 Outlier Rejection
IV Results and Discussion
IV.1 Axis Errors
@@
We also analyzed our idealized mock data set, finding an error of $2.3^{\circ}$, in line with the findings of Häussler et al. (2007) for idealized galaxies. Given the idealized nature of this data, we would expect that features like bars, warps, and spirals would produce a larger uncertainty in a real observation. However, a good fit to the ideal data provides validation of the methods used for finding the position angle. The difference between the results from Illustris and idealized data gives an indication of the impact of nonideal features, such as bars or warps, on the observed position angle.
Figure 7: Axis ratios calculated from Illustris. An overall uncertainty of 0.087 was found, with similar values across inclinations of $20^{\circ}$, $40^{\circ}$, and $60^{\circ}$. The dashed vertical line marks the median axis ratio, and the shaded region shows the central $68\%$ interval.
Turning our attention to the errors in the axis ratio, shown in Figure 7, we find the Illustris sample is well fit with a beta distribution across all inclination angles. An overall uncertainty of 0.087 was found over the entirety of the dataset. That value is improved to between $\sigma_{q}=0.078$ and $\sigma_{q}=0.085$ if we consider each inclination individually. However, that information would not be available in a real observation, so we opt to use the larger value in our calculations. Note, the axis ratio does not match what we would expect considering an inclined thin disk. Instead, all axis ratios trend greater than a theoretical value, with discrepancy increasing with inclination. We believe this to be an effect due to limitations in the ideal thin disk calculation, rather than an error in the fitting. The axis ratio in the calculation of Equation 38 depends on the true observed ratio, not on a thin disk approximation.
@@
With the uncertainty in the photometric axes established, the model is then characterized on an idealized mock data set, described in Section III.6.1. A set of fiducial values of $\gamma_{\times}=0$, $\phi=0^{\circ}$, $i=40^{\circ}$, and $\delta_{V}=$10\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$ were held constant across all runs, with the exception of the parameter being studied, which was varied across the range given in Table 2. A total of 50 realizations of the data are used for each parameter configuration.
IV.2 Idealized Mock Data
Figure 8: Characterization of the algorithm across the mock data set. A fiducial set of parameters: $\gamma_{\times}=0$, $\phi=0^{\circ}$, $i=40^{\circ}$, and $\delta_{V}=$10\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$, was used while the individual parameters were varied. In all but the top-left plot, the distribution of $\Delta\gamma_{\times}\equiv\gamma_{\times}^{\mathrm{fit}}-\gamma_{\times}^{\mathrm{true}}$ is equivalent to the distribution of fitted $\gamma_{\times}$ values for the fiducial case $\gamma_{\times}^{\mathrm{true}}=0$. Across each trial is 50 realizations of the same configuration. In the violin plots, orange lines mark the median of the data while the boxes give the interquartile range, and whiskers extend to 5th and 95th percentiles. The combined model is given in purple, kinematic-only in green, and a parametric kinematic model in blue.
@@ IV.2 Idealized Mock Data
We compared three approaches to fitting the data: (1) a nonparametric model using only kinematic data, (2) a nonparametric model using both kinematic and photometric data, and (3) a simple parametric model using only kinematic data based on Equation 40, henceforth referred to as kinematic, combined, and parametric, respectively. To quantify the performance of each, we examine the distribution of $\Delta\gamma_{\times}\equiv\gamma_{\times}^{\mathrm{fit}}-\gamma_{\times}^{\mathrm{true}}$, which reduces to the distribution of $\gamma_{\times}^{\mathrm{fit}}$ for $\gamma_{\times}^{\mathrm{true}}=0$. The results are shown in Figure 8.
@@
While varying the value of $\gamma_{\times}$ from 0.2-0.2 to $0.2$, we find the kinematic approach generally outperforms the parametric model, and the combined model improves on the kinematic only. At larger values of shear, there is a ${\sim}{3}\sigma$ underfitting bias in the combined model. This is a possible indication on the limits of the approximations used in deriving the theoretical difference between the kinematic and photometric position angles, particularly in the small angle approximation used in Equation 5. These systematic offsets could, in principle, be characterized with additional mock tests and used to define an empirical calibration relation for correcting the fitted shear values in future applications.
@@
The effects of inclination are most dramatic for the parametric model, with minimal ability to fit at the lowest inclination angle. The distribution of $\gamma_{\times}$ shows the parametric model running into the bounds on the parameter at $i=10^{\circ}$. The kinematic model performs better than the parametric at low inclination and improves with inclination up to $60^{\circ}$. The combined model performs well over a wide range of inclinations, with the smallest error at mid-inclinations from $40^{\circ}$$60^{\circ}$. Increasing to the highest inclination tested, $80^{\circ}$, all models have greater uncertainty than at mid-inclinations. Both nonparametric approaches give similar results at higher inclination.
@@
The information available from the kinematic and photometric data varies with inclination. The information available in the photometric data is greatest at low inclinations. The apparent axis ratio increases with decreasing inclination, causing Equation 11 to diverge as $q$ approaches unity and trend to zero as the galaxy approaches edge-on. The velocity field, however, is featureless at low inclinations. The line-of-sight velocity is a function of $\sin i$, with the differential velocity increasing with inclination. As the galaxy approaches edge-on, the minor axis vanishes in an infinitely thin disk model, causing the increase in uncertainty we see at $i=80^{\circ}$.
In the top-right panel of Figure 8, we explore the response of each model to a change in the intrinsic (source-plane) position angle of the galaxy. The fiducial value of shear is $\lvert\gamma\rvert=\gamma_{\times}=0$, so all of the effects are due to behavior of the fitting algorithm in response to position angle, and not mixing of shear components. To explore this behavior, we do not make any attempt to align the galaxy axes with the coordinate axes. The fit for $\gamma_{\times}$ shows dramatic effects as the position angle approaches $45^{\circ}$. The kinematic-only model is completely insensitive to $\gamma_{\times}$ at $\phi=45^{\circ}$. The shear transformation back to the source plane happens before the position-angle rotation and, with the galaxy at $45^{\circ}$, $\gamma_{\times}$ (applied with reference to the coordinate axes as opposed to the galaxy axes) will behave as $\gamma_{+}$ in the galaxy frame and only produce stretching and compressing along the galaxy axes and have no effect on galaxy symmetry. This can be seen explicitly in Equation 22. For the case of $\phi=45^{\circ}$,
$\mathcal{M}_{\text{major}}$ $=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$, $\mathcal{M}_{\text{minor}}$ $=\begin{pmatrix}0&-1\\ -1&0\end{pmatrix}$, (43)
completely independent of $\gamma_{\times}$.
@@
The combined model, on the other hand, seems to strongly outperform all other models at that position angle. This is not the result of a good fit, but rather is due to the fitting process giving $\gamma_{\times}=0$ for any value of true shear. If we revisit Equation 17, we can see that at $\phi=45^{\circ}$, the detector-frame kinematic position angle will also be $45^{\circ}$. This will push $\gamma_{\times}$ to zero in Equation 9, regardless of true shear. This is the greatest difference in the two different nonparametric approaches, with both approaches failing for different reasons. However, in practice, the fitting can be done with the observed kinematic major axis aligned with the coordinate axis prior to fitting to minimize these effects.
@@
Finally, we look at the effect of noise added to the kinematic data. Gaussian noise is added at 1, 5, 10, 20, and $40\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$ on the central pixel and scales with the inverse of the square root of the local intensity, given by Equation 41. All models show trends of a better fit at low noise and poor fit at high noise. The combined approach increasingly outperforms the other methods as noise increases. The quality of the parametric fit at extremely low noise is probably an artifact of the mock data and parametric model using the same functional form and would not translate into performance in a real observation. At $40\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$, the distribution of $\gamma_{\times}$ shows the parametric model is running into the bounding limits and is almost completely unable to fit for shear.
IV.2.1 Effect of $\gamma_{+}$
Figure 9: Impact of $\gamma_{+}$ on the fit for $\gamma_{\times}$ using the combined model on idealized mock data. The same fiducial set of parameters: $\gamma_{\times}=0$, $\phi=0^{\circ}$, $i=40^{\circ}$, and $\delta_{V}=$10\text{\,}\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$, was used as in Figure 8. In the top-left panel, the total magnitude of shear depends on the values of both $\gamma_{\times}$ and $\gamma_{+}$, with a maximum $\lvert\gamma\rvert=0.25$. For all other panels, $\gamma_{\times}=0$, and thus $\lvert\gamma\rvert=\lvert\gamma_{+}\rvert$. The top-right panel shows the effects of galaxy position angle, which effectively mixes $\gamma_{\times}^{\mathrm{gal}}$ and $\gamma_{+}^{\mathrm{gal}}$ components. While $\gamma_{\times}=0$ in the fixed detector coordinate system, it is nonzero in a coordinate system aligned with the galaxy axes. For nonzero values of $\gamma_{+}$ and $\phi$, the actual value being fit is $\gamma_{\times}^{\mathrm{fit}}\simeq\gamma_{\times}^{\mathrm{true}}-\gamma_{+}^{\mathrm{true}}\tan 2\phi$, which is plotted in red (with the exception of at $\phi=45^{\circ}$, where the term is undefined). The bottom-left panel shows the effects of inclination angle. At low inclination, $\gamma_{+}$ causes the photometric axis to be vertical, which contributes to the observed bimodal distributions. The bottom-right panel shows the effect of $\gamma_{+}$ on the fiducial setup.
@@ IV.2.1 Effect of $\gamma_{+}$
While the MIRoRS algorithm focuses primarily on solving for $\gamma_{\times}$, real observations will contain both components of shear. We explore the impact of $\gamma_{+}$ on our fit for $\gamma_{\times}$ using the combined model, using idealized mock data. We show the distribution of $\Delta\gamma_{\times}$ and scatter in the fit for $\gamma_{\times}$ in Figure 9. In all but the top-left panel, $\Delta\gamma_{\times}$ reduces to $\gamma_{\times}^{\mathrm{fit}}$. We do not attempt to fit for $\gamma_{+}$, but believe our methods could be improved by considering the parameter, which we leave for future work.
Overall, the effect of positive values of $\gamma_{+}$ is to elongate the galaxy along the $y$-axis, the minor axis of our typical galaxy orientation. For small values of $\gamma_{+}$, this manifests primarily as an error in the axis ratio used in fitting for the observed kinematic position angle. The top-left panel of Figure 9 shows the effect of varied $\gamma_{\times}$ at each fixed value of $\gamma_{+}$. The total magnitude of shear depends on both components with $\lvert\gamma\rvert=\sqrt{\gamma_{+}^{2}+\gamma_{\times}^{2}}$ and reaches a maximum value of $0.25$ in the case of $\gamma_{+}=0.15$ and $\gamma_{\times}=\pm 0.2$. The addition of $\gamma_{+}$ increases scatter, particularly at low $\gamma_{\times}$ values. At larger values of $\gamma_{\times}$, the uncertainty is comparable across all tested values of $\gamma_{+}$. The bias exhibited in Figure 8 is also present. We can also see this increase in scatter in the bottom-right panel, where we look explicitly at the fiducial model while varying only $\gamma_{+}$. For $\gamma_{+}\leq 0.15$, there is no bias in the fit for $\gamma_{\times}$, however at $\gamma_{+}=0.2$, the fit is biased and actually bimodal.
This bimodality is evidence of a distinct failure-mode of the MIRoRS method with nonzero $\gamma_{+}$. The elongation of the unlensed minor axis by $\gamma_{+}$ switches the photometric major and minor axes at sufficiently large values of $\gamma_{+}$ or low inclination angles (bottom-left panel of Figure 9), causing the observed photometric major axis to align with the kinematic minor. This is a blatant violation of our assumption that the photometric transformation can be treated as a simple rotation. This produces an increased scatter as the observed $\pi/2$ difference between kinematic and photometric position angles causes the photometric term to bias the fit toward large values of shear via Equation 13. The bimodality arises as the sign of the photometric position angle is ambiguous at $\pm\pi/2$. As large axis ratios and large values of $\gamma_{\times}$ are generally incompatible, this effect can be corrected by identifying the instances where shear has introduced the $\pi/2$ angle offset and defining an effective photometric position angle corresponding to the unlensed major axis. While this poses a problem that must be addressed to realize the full benefits of including the photometric position angle, $\gamma_{+}$ (at $\phi=0^{\circ}$) does not affect the symmetry of the velocity field.
The top-right panel of Figure 9 shows how the source-plane position angle affects the fit for $\gamma_{\times}$. In the detector coordinates, a fiducial value of $\gamma_{\times}=0$ is applied with varied values for $\gamma_{+}$. As the position angle approaches $45^{\circ}$, $\gamma_{+}$ applied in the fixed coordinates acts as $\gamma_{\times}$ in a coordinate system aligned with the galaxy axes. The MIRoRS algorithm, as stated in Equation 22, first applies inverse shear to bring the observed galaxy back to the intrinsic source plane and then applies rotation by $\phi$ to align the galaxy axes with the coordinate axes before the reflection operations. Thus, the fitted components of shear are defined in the fixed detector frame, and $\phi$ alone is unable to compensate for changes in shear orientation.
Ultimately, the optimizer is attempting to compensate for the symmetry-breaking shear component in the galaxy frame with $\gamma_{\times}$ applied in the detector frame. For model data created with Equation 1, the full transformation matrix to first order in shear, including rotations, is
$\left(\begin{smallmatrix}\mathrm{c}2\phi&\mathrm{s}2\phi-2\gamma_{+}\mathrm{s}2\phi+2\gamma_{\times}\mathrm{c}2\phi\\ \mathrm{s}2\phi+2\gamma_{+}\mathrm{s}2\phi-2\gamma_{\times}\mathrm{c}2\phi&-\mathrm{c}2\phi\end{smallmatrix}\right),$ (44)
where s and c represent sine and cosine operations. In contrast, the fitting algorithm, without access to $\gamma_{+}$, is fitting with
$\left(\begin{smallmatrix}\mathrm{c}2\phi&\mathrm{s}2\phi+2\gamma_{\times}\mathrm{c}2\phi\\ \mathrm{s}2\phi-2\gamma_{\times}\mathrm{c}2\phi&-\mathrm{c}2\phi\end{smallmatrix}\right).$ (45)
Our optimizer effectively seeks to minimize the difference between the two transformations and, from the cross-terms, produces
$\gamma_{+}^{\mathrm{true}}\sin 2\phi+\gamma_{\times}^{\mathrm{true}}\cos 2\phi-\gamma_{\times}^{\mathrm{fit}}\cos 2\phi\simeq 0.$ (46)
The actual quantity being fit is then
$\gamma_{\times}^{\mathrm{fit}}\simeq\gamma_{\times}^{\mathrm{true}}-\gamma_{+}^{\mathrm{true}}\tan 2\phi,$ (47)
where $\gamma_{\times}^{\mathrm{true}}$ and $\gamma_{+}^{\mathrm{true}}$ are the true detector-frame values. This quantity, with $\gamma_{\times}^{\mathrm{true}}=0$, is shown in the top-right panel of Figure 9 as red lines. While the MIRoRS algorithm fits for this theoretical value well, care should be taken to recover $\gamma_{\times}$ independent of $\gamma_{+}$ and $\phi$.
To fit for the true cross-component of shear, the galaxy must be aligned such that $\phi=0$ (assuming $\gamma_{+}$ is nonzero). Assuming equal magnitudes of $\gamma_{+}$ and $\gamma_{\times}$, even small values of $\phi$ can bias the fit ($5^{\circ}$ would produce a bias of ${\approx}17\%$). However, $\phi$ is the source-plane position angle and is not directly observable. Notably, aligning the coordinates with the observed photometric axes can introduce significant bias, given by Equation 8. Aligning with the observed kinematic axes produces a much smaller bias with $\phi\approx\gamma_{\times}$.
Figure 10: Distribution of $\gamma_{\times}$ for mock data constructed with nonzero $\gamma_{+}$ and $\phi$ after an initial rotation to minimize $\phi$ and capture shear in the galaxy coordinate frame. The true galaxy frame $\gamma_{\times}$ is shown in red.
To explore these effects, we apply an initial rotation to align the detector coordinate $x$-axis with the observed kinematic axis and fit with the standard MIRoRS algorithm. We fit for the kinematic major axis by maximizing a directional gradient, but admit more sophisticated methods are likely better (see Krajnovic et al. (2006)). Adding an additional fitting parameter for the initial rotation is largely degenerate with the rotation in the source plane, complicating the fit. However, it is likely possible that an iterative process could be used that starts at the observed kinematic position angle and applies a rotation according to Equation 11 until the fitted value of $\phi$ is pushed to zero, which we leave for future work. The resulting fit is shown in Figure 10. For most values of $\gamma_{+}$ and $\phi$, the fitted value matches the expected value of $\gamma_{\times}$ in the galaxy reference frame. At the largest values of applied shear, there is an underfitting, matching what is was seen in the top-left panels of Figures 8 and 9.
@@
IV.2.2 Effect of Asymmetries
Figure 11: Example of a mock velocity field with an imposed linear drift in kinematic position angle used to test sensitivity to spatially correlated structure. The curved line marks the kinematic major axis as a function of radius for a variation of $35^{\circ}$ across the field.
To explore the effects of spatially correlated asymmetries, we created mock velocity fields with a kinematic position angle that varies linearly with radius, $\mathrm{PA}(r)=\phi+\mathrm{PA}_{k}\,r$, where $\mathrm{PA}_{k}$ is the imposed PA-variation rate. This produces a velocity field (see Figure 11) that violates both reflection symmetries in a coherent manner, and is the most common type of asymmetry found across a survey of 2776 MaNGA galaxies (Stark et al., 2018). We generated 50 realizations for variation rates of $\pm 5^{\circ}$, $\pm 15^{\circ}$, $\pm 25^{\circ}$, and $\pm 35^{\circ}$ across the field using the same fiducial parameters described in §IV.2. The results shown in Figure 12 exhibit scatter comparable to the symmetric case, but with a systematic bias that increases with $\mathrm{PA}_{k}$. The largest variation rate produced a median recovered shear of $\gamma_{\times}=0.0207$, a $3.5\sigma$ deviation of the median, where $\sigma$ is the standard error of the median.
We then applied the Moran’s $I$ test (bottom panel of Figure 12) and found that the statistic increases monotonically with $\lvert{\mathrm{PA}_{k}}\rvert$, as expected for increasingly coherent departures from symmetry. Applying a selection cut of $p\leq 0.001$ rejected $\sim 70\%$ of realizations at the highest variation rates while rejecting only $\sim 15\%$ in the fiducial case, confirming the efficacy of the Moran test in identifying asymmetric or warped fields. Quantifying these effects also opens a path toward correcting them (e.g. by explicitly estimating the degree of asymmetry with a Radon transform (Stark et al., 2018)). In any sufficiently large ensemble, these orientation-dependent effects are expected to average away.
Figure 12: Characterization of the MIRoRS algorithm on mock galaxies with a linearly varying kinematic position angle. Top panel: fitted scatter across different imposed PA-variation rates. The middle panel shows the distribution of fitted $\gamma_{\times}$ values and corresponding bias, which remains small but nonzero due to the coherent PA drift. The bottom panel shows Moran’s II (blue violins) and the fraction of fits rejected at the $p\leq 10^{-3}$ level (red dashed curve).
IV.3 Illustris Data
Figure 13: Comparison of fitted $\gamma_{\times}$ values for all Illustris halos at inclinations of $20^{\circ}$, $40^{\circ}$, and $60^{\circ}$. Top panel: uncertainty in $\gamma_{\times}$ for each method as a function of inclination, comparing the sample filtered with Moran’s $I$ test (red), the combined model (purple), the kinematic-only model (green), and the parametric model of DW23 (blue). Bottom panel: histograms showing the distribution of fitted $\gamma_{\times}$ in each inclination bin. Each model is displayed as a row with the same color-coding as above. The preferred two-component GMM fit is marked by a solid line with the core (broad) component as a dashed (dotted) line. All histograms are plotted with the same scaling for equal comparison. The true shear for all halos is $\gamma_{\times}=0$.
Extending our analysis to the Illustris data, we fit the 358 halos at three inclinations each ($20^{\circ}$, $40^{\circ}$, and $60^{\circ}$) with both the kinematic and combined models. These halos are from a z=0 snapshot and are all expected to have $\gamma_{\times}=0$. Results are shown in Table 3, along with a comparison to analysis of the same data using a parametric model by DW23 and results after filtering with the Moran’s $I$ test. We report the overall uncertainty in the distribution of fitted $\gamma_{\times}$ values, but find that the majority of the results are best fit with a two-component Gaussian mixture model (GMM) according to the Bayesian information criteria. We report the scatter in the narrow component of the GMM separately as $\sigma_{\mathrm{core}}$. We show the distribution of results along with the overall uncertainty for each case in Figure 13.